Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 123     Entry time: 09 Jun 2024, 22:45
Author: Chan Kim 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 7.0 keV, up to 202 pulses at 564 kHz, 1.0 mJ.

Optical laser delivery

  • ns laser for jet illumination

Achievements/Observations

  • Data collection for user sample through column

Issues

  • Pressure spike while rotating nozzle tripped AGIPD
  • Temperature of ASICs of Q2M1 module too low and not easy to turn it on (issue solved after changing the bath temperature to -27 degreeC)
  • Julabo chiller issue again (In the end, set temperature changed to -29.5 degreeC and continued the experiment)
  • Another 2 ASICs in Q2M2 and Q2M3 seem unhappy
ELOG V3.1.4-7c3fd00